If the width of a certain rectangle is increased by 12 and length reduced by 36 we get a square with the same area as original rectangle. Find length and width of orignal rectangle?
Answer by hamsanash1981@gmail.com(151)

Let the side of the square formed be x, then lenght = x+36, Breath = x-12
given the area of rectangle = area of the square.
(x+36)(x-12) = X^2
X^2 +36x-12X+432 = x^2
24x+432=0
24x = 432
x = 18
Length = 18+36 = 54
Breath = 18 - 12 = 6